# Strong Quantum Mpemba Effect with Squeezed Thermal Reservoirs

### v1

@@ 3 QMpE with squeezed thermal baths
The system of interest in this work is a single quantum two-level system interacting with a squeezed thermal bath, at temperature $T$. Such a kind of reservoir is efficiently engineered in the current state-of-the-art in nanobeam experiments with squeezed electronic noise [22], or in optical cavity systems [32, 33, 34]. As a model, such a system is described by the total Hamiltonian $\hat{H}_{T}=\hat{H}_{\mathrm{tls}}+\hat{H}_{\mathrm{b}}+\hat{H}_{\mathrm{int}}$. We assume that the two-level system has generic excited and ground states, respectively denoted by where $|e\rangle$ and $|g\rangle$, such that $\hat{H}_{\mathrm{tls}}=\hbar\omega_{0}(|e\rangle\langle e|-|g\rangle\langle g|)$. The thermal bath is described by a system of non-interacting quantum Harmonic oscillators as $\hat{H}_{\mathrm{b}}=\hbar\sum\nolimits_{k}\omega_{k}\hat{b}_{k}^{\dagger}\hat{b}_{k}$, and the bath interacts with the two-level system through the interaction Hamiltonian
@@
where $\hat{\sigma}^{+}=(\sigma^{-})^{\dagger}=|g\rangle\langle e|$, and $g_{k}$ is the coupling of the atom with the bath mode with wave vector $k$. The reservoir temperature and squeezing properties are introduced by assuming each mode of the reservoir initially at the state $\hat{\rho}_{b,k}(0)=\hat{S}_{k}(r,\varphi)\hat{\rho}_{b,k}^{\mathrm{th}}(0)\hat{S}_{k}^{\dagger}(r,\varphi)$, where $\hat{\rho}_{b,k}^{\mathrm{th}}(0)$ is the thermal state of the bath modes at temperature $T$, and $\hat{S}_{k}(r,\varphi)=\exp\big{[}\frac{1}{2}\xi_{k}^{\ast}\hat{b}_{k}^{2}-\mathrm{h.c}\big{]}$ is the squeezing operator, with the squeezing parameters $r_{k}$ and $\varphi_{k}$ encoded in $\xi_{k}=r_{k}e^{i\varphi_{k}}$ [35]. For simplicity, we assume $\xi_{k}=\xi=re^{i\varphi}$ for all $k$. Under these conditions, the dynamics for the two-level system is governed by Eq. (2), with the effective Lindblad superoperator $\mathfrak{L}[\bullet]$ as
@@
with $\hat{R}_{1}=\sqrt{\gamma_{0}(N_{\mathrm{th}}+1)}\hat{R}(r,\varphi)$, and $\hat{R}_{2}=\sqrt{\gamma_{0}N_{\mathrm{th}}}\hat{R}^{\dagger}(r,\varphi)$, where $\hat{R}(r,\varphi)=\cosh(r)\hat{\sigma}^{-}+e^{i\varphi}\sinh(r)\hat{\sigma}^{+}$, $\hat{\sigma}^{-}=\hat{\sigma}^{+\dagger}=|g\rangle\langle e|$, $\gamma_{0}$ the spontaneous emission rate, and we assume the Bose-Einstein statistics for the average number of thermal photons $N_{\mathrm{th}}=1/(e^{\hbar\omega_{0}/k_{B}T}-1)$. The above dynamics is valid under the Born-Markov and rotating wave approximations [36, 37].
In this framework, the hot and cold initial states $\hat{\rho}_{\mathrm{h}}$ and $\hat{\rho}_{\mathrm{c}}$ are prepared independently as the steady-state of the Eq. (7) for different temperatures $T_{\mathrm{h}}$ and $T_{\mathrm{c}}$, respectively, for a given predefined set of squeezed parameters $\{r,\varphi\}$. In both cases, the Hamiltonian for the two-level system is the same, and we consider it as given by the Landau-Zener
@@
which can be implemented by a weak classical oscillating drive ($\Omega\ll\omega_{0}$) with frequency $\omega_{\mathrm{dr}}$ and phase $\phi$, and detuned of $\Delta_{0}=\omega_{\mathrm{dr}}-\omega_{0}$ from the atomic transition $\omega_{0}$. After this stage, each system is placed in contact with the reservoir at very low temperature $Tk_{B}\ll\hbar\omega_{0}$, but keeping the same squeezing parameters used to prepare the hot and cold states, such the system is driven from their respective initial states according to the Eq. (7).
Figure 3: Synchronized behavior for the parameter $\mathcal{M}$, the ratio between the populations of the coldest reservoir modes in the hot $I_{B}^{\mathrm{hot}}$ and cold $I_{B}^{\mathrm{cold}}$ initial states, and the ratio $I_{B}^{\lambda}$ between the highest and lowest relaxation decay as function of the squeezing parameter $r$. The quantities $I_{B}^{\mathrm{hot}}$, $I_{B}^{\mathrm{cold}}$, and $I_{B}^{\lambda}$ are measured in units of Bel.
@@
In Fig. 2 we show the behavior of the Mpemba parameter $\mathcal{M}$ as function of the the squeezing parameter $r$ and the phase $\varphi$, for two different choices of ratio $\Omega/\Delta_{0}$, and the behavior of the quantities $\mathcal{B}_{\mathrm{c}}(t)$ and $\mathcal{B}_{\mathrm{h}}(t)$ along the evolution for each initial state. As conclusion, it is possible to conclude that the QMpE is not a common effect in the system, as it is only possible to observe it for a small range of parameters. In particular, we highlight the case in which $r=0.98$ and $\phi/\pi=0.025$, where the parameter $\mathcal{M}$ captures moderate and strong Mpemba effects in Fig. 2a and 2b, respectively. In fact, the Figs. 2c2d and 2e corresponds, respectively, to the points I, II and III of Fig. 2a, where one clearly observes that the parameter $\mathcal{M}$ captures how strong the QMpE is for each case. Furthermore, of particular interest we highlight the panels in Figs. 2f2g and 2h, related to the same points in Fig. 2b. By focusing on the cases II and III we can see that the QMpE gets stronger than in Fig. 2a, and it is mainly related to a faster decay of the hot state in both cases, while the cold states decay slower, leading then to a strong QMpE [15].
@@
While the results aforementioned illustrate the role of the parameter $\mathcal{M}$ introduced in this work, it does not explain anything about the QMpE. To this end, we now discuss how the conditions established previously allows us to understand the mechanism behind QMpE. To this end, we conveniently define some quantities to support our analysis. In this case, we know the system admits four eigenvalues $\lambda_{n}$, and then we define the quantity $I_{B}^{\lambda}=\log_{10}(\mathrm{Re}[\lambda_{3}]/\mathrm{Re}[\lambda_{1}])$, which provide the ratio between the real part of the fastest decay mode $\lambda_{3}$ and the slowest mode $\lambda_{1}$ in a Bel scale. This quantity is relevant to our analysis because $I_{B}^{\lambda}$ quantifies how many orders of magnitude bigger $\mathrm{Re}[\lambda_{3}]$ is compared with $\mathrm{Re}[\lambda_{1}]$. Concomitantly, we claim that the QMpE can only be fully understood if we also investigate the relations between the populations $\gamma_{n}$ for hot and cold states. Therefore, similarly, we also compute the same quantity defined for the “populations” $\gamma_{n}$ of the hot and cold states. Of interest to our analysis is the ratio between $I_{B}^{\mathrm{hot}}=\log_{10}(\mathrm{Re}[\gamma^{\mathrm{hot}}_{3}]/\mathrm{Re}[\gamma^{\mathrm{hot}}_{1}])$ and $I_{B}^{\mathrm{cold}}=\log_{10}(\mathrm{Re}[\gamma^{\mathrm{cold}}_{3}]/\mathrm{Re}[\gamma^{\mathrm{cold}}_{1}])$. Therefore, we can now explain the emergence of strong Mpemba effect in the system.
@@
In Fig. 2 we consider the set of parameters corresponding to the points III of the Figs. 2a and 2b. First we show, for each case, how $\mathcal{M}$ changes with respect to the squeezing parameter $r$, and we observe that the highest values of $\mathcal{M}$ in each example is achieved for different choices of $r$. This is the ideal scenario to test our conditions, so we investigate this behavior taking into account the Fig. 3. By increasing the squeezing parameter $r$, we create more squeezed thermal photons in the system, which directly affects the ratio $I_{B}^{\lambda}$. In principle, one could expect emergence of QMpE in for any $r>0.7$, because we have QMpE for $r=0.7$ and the decay rate of the mode $\lambda_{3}$ becomes even faster as $r$ increases. However, this analysis fails in our system because the slowest decaying mode is significantly populated with respect to the other modes. Therefore we invoke the additional conditions over $|\gamma^{\mathrm{hot}}_{n}|$ to explain this result.
@@
In Figs. 3a3b and 3c we observe that even when $|\gamma_{3,2}|$ is around one order of magnitude bigger than $|\gamma_{1}|$, the Mpemba parameter detect a moderate QMpE because the population in the mode $\lambda_{1}$ is big enough to suppress QMpE. However, by combining the fact that $|\gamma_{3,2}|\gg|\gamma_{1}|>0$, as seen in Fig. 3b, and $I_{B}^{\lambda}\sim 1$ ($\mathrm{Re}[\lambda_{3}]\sim 10\mathrm{Re}[\lambda_{1}]$), in Fig. 3c, we observe the maximum QMpE for the regime of parameters considered is achieved when $|\gamma_{3,2}|\gg|\gamma_{1}|$. In addition, this effect can be even stronger (strong QMpE) if we properly choose the Hamiltonian parameters from $\phi=0.025\pi$ to $\phi=0.075\pi$ (Figs. 3a3b and 3c), with $\mathcal{M}$ reaching values close to $0.75$. Similarly, our analysis explains the absence of QMpE even for the squeezed degree providing decay rates satisfying $\mathrm{Re}[\lambda_{3}]\gg\mathrm{Re}[\lambda_{1}]$. This result is relevant to our discussion because the Hamiltonian parameters can be efficiently controlled through external classical drives. In particular, we notice that in the regime in which QMpE is suppressed, such suppression is related to the fact that the hot state has a significant amount of population in the mode $\lambda_{1}$. When such a population is sufficiently reduced, we can achieve strong QMpE with $\mathcal{M}\approx 0.75$. It is worth to pay attention to the fact that, in all cases considered in Fig. 3, $I_{B}^{\mathrm{cold}}$ is small compared to $I_{B}^{\mathrm{hot}}$, which reinforces our claim that the conditions established are necessary to achieve strong QMpE.
4 Conclusions and prospects
@@
In conclusion, the results of our work are three-fold. First, we introduced a new parameter to efficiently witness the emergence of QMpE and to quantify how strong the QMpE is. Using this metric as figure of merit, we developed conditions able to predict scenarios where strong QMpE can be observed. Also, we showed how our results can be applied to explain the experimental observation of strong QMpE reported in Ref. [15]. To end, we also showed how to achieve different regimes of QMpE through of a single two-level system interacting with a squeezed thermal reservoir.
@@
We state that the QMpE superoperator theory and conditions developed in this study goes beyond the example considered in this study. For example, our results shade light over the experimental results reported in Ref. [15]. In this work the authors provided experimental data for the eigenmodes of the Lindbladian of a three-level system encoded in the electronic states of a rapped 40Ca+ ion. By driving such a system through the natural decay relaxation of the system, they investigate the behaviors of the populations of the cold and hot states with respect to the eigenmodes of $\mathfrak{L}[\bullet]$ (slightly different from our definition). As conclusion, they engineered an initial (pure) strong QMpE state $|sME\rangle$ able to induce a very fast decay of the hot state in comparison with the cold one. As a characteristic of such a state, through the experimental data, they observed that the component of $|sME\rangle$ over the slowest decaying state is numerically zero. This corresponds to the discussions done for the results shown in Fig. 3 of our work.
@@
It is worth to mention that, as aforementioned, our results can be experimentally reproduced in any experiment where squeezed thermal baths can be created or simulated. In particular, simulations of our findings can be done by digitizing the master equation considered in this work through squeezed Generalized Amplitude Damping channels [38]. In this regard, our results may be an sufficient, but not necessary, guide to describe both the reservoir and state engineering required to exploit new degrees of QMpE beyond the strong regime.

### v2

@@ III QMpE with squeezed thermal baths
The system of interest in this work is a single quantum two-level system interacting with a squeezed thermal bath at temperature $T$. As a model, such a system is described by the total Hamiltonian $\hat{H}_{T}=\hat{H}_{\mathrm{tls}}+\hat{H}_{\mathrm{b}}+\hat{H}_{\mathrm{int}}$. We assume that the two-level system has generic excited and ground states, respectively denoted as $|e\rangle$ and $|g\rangle$, such that $\hat{H}_{\mathrm{tls}}=\hbar\omega_{0}(|e\rangle\langle e|-|g\rangle\langle g|)$. The thermal bath is described by a system of non-interacting quantum harmonic oscillators as $\hat{H}_{\mathrm{b}}=\hbar\sum\nolimits_{k}\omega_{k}\hat{b}_{k}^{\dagger}\hat{b}_{k}$, and the bath interacts with the two-level system through the interaction Hamiltonian.
@@
where $\hat{\sigma}^{+}=(\sigma^{-})^{\dagger}=|g\rangle\langle e|$, and $g_{k}$ is the coupling of the atom with the bath mode with wave vector $k$. The reservoir temperature and squeezing properties are introduced by assuming each mode of the reservoir is initially in the state $\hat{\rho}_{b,k}(0)=\hat{S}_{k}(r,\varphi)\hat{\rho}_{b,k}^{\mathrm{th}}(0)\hat{S}_{k}^{\dagger}(r,\varphi)$, where $\hat{\rho}_{b,k}^{\mathrm{th}}(0)$ is the thermal state of the bath modes at temperature $T$, and $\hat{S}_{k}(r,\varphi)=\exp\big{[}\frac{1}{2}\xi_{k}^{\ast}\hat{b}_{k}^{2}-\mathrm{h.c}\big{]}$ is the squeezing operator, with the squeezing parameters $r_{k}$ and $\varphi_{k}$ encoded in $\xi_{k}=r_{k}e^{i\varphi_{k}}$ Manzano:18 . For simplicity, we assume $\xi_{k}=\xi=re^{i\varphi}$ for all $k$. Under these conditions, the dynamics for the two-level system is governed by Eq. (2), with the effective Lindblad superoperator $\mathfrak{L}[\bullet]$ as.
@@
with $\hat{R}_{1}=\sqrt{\gamma_{0}(N_{\mathrm{th}}+1)}\hat{R}(r,\varphi)$, and $\hat{R}_{2}=\sqrt{\gamma_{0}N_{\mathrm{th}}}\hat{R}^{\dagger}(r,\varphi)$, where $\hat{R}(r,\varphi)=\cosh(r)\hat{\sigma}^{-}+e^{i\varphi}\sinh(r)\hat{\sigma}^{+}$, $\hat{\sigma}^{-}=\hat{\sigma}^{+\dagger}=|g\rangle\langle e|$, $\gamma_{0}$ is the spontaneous emission rate, and we assume that the Bose-Einstein statistics applies for the average number of thermal photons $N_{\mathrm{th}}=1/(e^{\hbar\omega_{0}/k_{B}T}-1)$. The above dynamics is valid under the Born-Markov and rotating wave approximations Dalton:99 ; Banerjee:10 .
This type of reservoir has been efficiently engineered in state-of-the-art nanobeam experiments using squeezed electronic noise Klaers:17 , where squeezing parameters $r$ tunable from 0 up to approximately 1 have been demonstrated. For this reason, we will consider this regime of squeezing in our analysis. Comparable levels of squeezing control have also been achieved in optical cavity systems Ourjoumtsev:11 ; Vahlbruch:16 and in parametric up-conversion and down-conversion experiments Vahlbruch:08 ; Heinze:22 , making these platforms promising candidates for the observation of the quantum Mpemba effect.
In this framework, the hot and cold initial states $\hat{\rho}_{\mathrm{h}}$ and $\hat{\rho}_{\mathrm{c}}$ are prepared independently as the steady-state of the Eq. (7) for different temperatures $T_{\mathrm{h}}$ and $T_{\mathrm{c}}$, respectively, for a given predefined set of squeezed parameters $\{r,\varphi\}$. In both cases, the Hamiltonian for the two-level system is the same, and we consider it as being given by the Landau-Zener
@@
which can be implemented by a weak classical oscillating drive ($\Omega\ll\omega_{0}$) with frequency $\omega_{\mathrm{dr}}$ and phase $\phi$, and detuned by $\Delta_{0}=\omega_{\mathrm{dr}}-\omega_{0}$ from the atomic transition $\omega_{0}$. After this stage, each system is placed in contact with the reservoir at very low temperature $Tk_{B}\ll\hbar\omega_{0}$, but keeping the same squeezing parameters used to prepare the hot and cold states, so that the system is driven from their respective initial states according to the Eq. (7).
Figure 3: Synchronized behavior of key quantities associated with the quantum Mpemba effect (QMpE) as a function of the squeezing parameter $r$, for two fixed values of the squeezing phase: $\varphi/\pi=0.025$ (left panels) and $\varphi/\pi=0.075$ (right panels). (a, d) Mpemba parameter $\mathcal{M}$, quantifying the presence of QMpE, shows distinct peaks indicating optimal conditions for the effect as $r$ is varied. (b, e) Corresponding logarithmic ratios (in bels) of the populations of the coldest reservoir modes for the hot ($I_{B}^{\mathrm{hot}}$) and cold ($I_{B}^{\mathrm{cold}}$) initial states. (c, f) Logarithmic ratio $I_{B}^{\lambda}$, representing the spread between the highest and lowest relaxation decay rates (in bels). The panels share a common x-axis (squeezing parameter $r$), allowing for direct comparison and revealing a synchronized behavior among $\mathcal{M}$, $I_{B}^{\mathrm{hot}}$, $I_{B}^{\mathrm{cold}}$, and $I_{B}^{\lambda}$. This synchronization highlights how the emergence of QMpE is linked to changes in both the reservoir mode populations and the decay rate spectrum.
@@
In Fig. 2 we show the behavior of the Mpemba parameter $\mathcal{M}$ as a function of the squeezing parameter $r$ and the phase $\varphi$, for two different choices of ratio $\Omega/\Delta_{0}$, and the behavior of the quantities $\mathcal{B}_{\mathrm{c}}(t)$ and $\mathcal{B}_{\mathrm{h}}(t)$ along the evolution for each initial state. In conclusion, it is possible to conclude that the QMpE is not a common effect in the system, as it is only possible to observe it for a small range of parameters. In particular, we highlight the case in which $r=0.98$ and $\phi/\pi=0.025$, where the parameter $\mathcal{M}$ captures strongest Mpemba effect observed in our system, for the regime of parameters shown in Fig. 2a and 2b, respectively. In fact, the Figs. 2c2d and 2e correspond to the points I, II, and III of Fig. 2a, where one clearly observes that the parameter $\mathcal{M}$ captures how strong the QMpE is for each case. Furthermore, of particular interest, we highlight the panels in Figs. 2f2g and 2h, related to the same points in Fig. 2b. By focusing on the cases II and III, we can see that the QMpE gets stronger than in Fig. 2a, and it is mainly related to a faster decay of the hot state in both cases, while the cold states decay slower, leading to a stronger QMpE Zhang:24 .
It means that our analysis reveals that the conditions we propose are more flexible than those required for the genuinely strong quantum Mpemba effect (QMpE) as defined in Ref.Zhang:24 , where the effect was experimentally demonstrated using a three-level system encoded in the low-lying electronic states of a single 40Ca+ ion. The strong QMpE necessitates a complete suppression of the population associated with the hot initial state in the subspace corresponding to the slowest-decaying eigenmode. This requirement is consistent with our framework under the constraint $0=|\gamma^{\mathrm{h}}{1}|\leq|\gamma^{\mathrm{h}}{2}|\leq\cdots\leq|\gamma^{\mathrm{h}}_{D^{2}-1}|$, which represents an extreme case of our general inequalities. Therefore, while our Mpemba parameter captures an enhanced form of QMpE, it remains a weaker manifestation compared to the stronger version observed in Ref.Zhang:24 .
@@
While the results aforementioned illustrate the role of the parameter $\mathcal{M}$ introduced in this work, it does not explain anything about the QMpE. To this end, we now discuss how the conditions established previously allow us to understand the mechanism behind QMpE. To this end, we conveniently define some quantities to support our analysis. In this case, we know the system admits four eigenvalues $\lambda_{n}$, and then we define the quantity $I_{B}^{\lambda}=\log_{10}(\mathrm{Re}[\lambda_{3}]/\mathrm{Re}[\lambda_{1}])$, which provides the ratio between the real part of the fastest decay mode $\lambda_{3}$ and the slowest mode $\lambda_{1}$ in a Bel scale. This quantity is relevant to our analysis because $I_{B}^{\lambda}$ quantifies how many orders of magnitude bigger $\mathrm{Re}[\lambda_{3}]$ is compared with $\mathrm{Re}[\lambda_{1}]$. Concomitantly, we claim that the QMpE can only be fully understood if we also investigate the relations between the populations $\gamma_{n}$ for hot and cold states. Therefore, similarly, we also compute the same quantity defined for the “populations” $\gamma_{n}$ of the hot and cold states. Of interest to our analysis is the ratio between $I_{B}^{\mathrm{hot}}=\log_{10}(\mathrm{Re}[\gamma^{\mathrm{hot}}_{3}]/\mathrm{Re}[\gamma^{\mathrm{hot}}_{1}])$ and $I_{B}^{\mathrm{cold}}=\log_{10}(\mathrm{Re}[\gamma^{\mathrm{cold}}_{3}]/\mathrm{Re}[\gamma^{\mathrm{cold}}_{1}])$. Therefore, we can now explain the emergence of different regimes of Mpemba effect in the system.
@@
In Fig. 2 we consider the set of parameters corresponding to the points III of the Figs. 2a and 2b. First we show, for each case, how $\mathcal{M}$ changes with respect to the squeezing parameter $r$, and we observe that the highest values of $\mathcal{M}$ in each example are achieved for different choices of $r$. This is the ideal scenario to test our conditions, so we investigate this behavior taking into account the Fig. 3. By increasing the squeezing parameter $r$, we create more squeezed thermal photons in the system, which directly affects the ratio $I_{B}^{\lambda}$. In principle, one could expect emergence of QMpE for any $r>0.7$, because we have QMpE for $r=0.7$ and the decay rate of the mode $\lambda_{3}$ becomes even faster as $r$ increases. However, this analysis fails in our system because the slowest decaying mode is significantly populated with respect to the other modes. Therefore we invoke the additional conditions over $|\gamma^{\mathrm{hot}}_{n}|$ to explain this result.
@@
In Figs. 3a3b and 3c we observe that even when $|\gamma_{3,2}|$ is around one order of magnitude bigger than $|\gamma_{1}|$, the Mpemba parameter detect a moderate QMpE because the population in the mode $\lambda_{1}$ is big enough to suppress QMpE. However, by combining the fact that $|\gamma_{3,2}|\gg|\gamma_{1}|>0$, as seen in Fig. 3b, and $I_{B}^{\lambda}\sim 1$ ($\mathrm{Re}[\lambda_{3}]\sim 10\mathrm{Re}[\lambda_{1}]$), in Fig. 3c, we observe the maximum QMpE for the regime of parameters considered is achieved when $|\gamma_{3,2}|\gg|\gamma_{1}|$. In addition, this effect can be even stronger if we properly choose the Hamiltonian parameters from $\phi=0.025\pi$ to $\phi=0.075\pi$ (Figs. 3a3b and 3c), with $\mathcal{M}$ reaching values close to $0.75$. Similarly, our analysis explains the absence of QMpE even for the squeezed degree providing decay rates satisfying $\mathrm{Re}[\lambda_{3}]\gg\mathrm{Re}[\lambda_{1}]$. This result is relevant to our discussion because the Hamiltonian parameters can be efficiently controlled through external classical drives. In particular, we notice that in the regime in which QMpE is suppressed, such suppression is related to the fact that the hot state has a significant amount of population in the mode $\lambda_{1}$. When such a population is sufficiently reduced, we can achieve stronger QMpE with $\mathcal{M}\approx 0.75$. It is worth paying attention to the fact that, in all cases considered in Fig. 3, $I_{B}^{\mathrm{cold}}$ is small compared to $I_{B}^{\mathrm{hot}}$, which reinforces our claim that the conditions established are necessary to achieve stronger QMpE.
IV Conclusions and prospects
@@
In conclusion, the results of our work are three-fold. First, we introduced a new parameter to efficiently witness the emergence of QMpE and to quantify how strong the QMpE is. Using this metric as a figure of merit, we developed conditions able to predict scenarios where different regimes of QMpE can be observed. Also, we showed how our results can be applied to explain the experimental observation of strong QMpE reported in Ref. Zhang:24 . Finally, we also showed how to achieve different regimes of QMpE through a single two-level system interacting with a squeezed thermal reservoir.
@@
We state that the QMpE superoperator theory and conditions developed in this study go beyond the example considered in this study. For example, our results shed light on the experimental results reported in Ref. Zhang:24 . In this work, the authors provided experimental data for the eigenmodes of the Lindbladian of a three-level system encoded in the electronic states of a trapped 40Ca+ ion. By driving such a system through the natural decay relaxation of the system, they investigate the behaviors of the populations of the cold and hot states with respect to the eigenmodes of $\mathfrak{L}[\bullet]$ (slightly different from our definition). As a conclusion, they engineered an initial (pure) strong QMpE state $|sME\rangle$ able to induce a very fast decay of the hot state in comparison with the cold one. As a characteristic of such a state, through the experimental data, they observed that the component of $|sME\rangle$ over the slowest decaying state is numerically zero. This corresponds to the discussions presented for the results shown in Fig. 3 of our work.
@@
It is worth mentioning that, as aforementioned, our results can be experimentally reproduced in any experiment where squeezed thermal baths can be created or simulated. In particular, simulations of our findings can be done by digitizing the master equation considered in this work through squeezed Generalized Amplitude Damping channels Srikanth:08 . In this regard, our results may be a sufficient, but not necessary, guide to describe both the reservoir and state engineering required to exploit new degrees of QMpE. Also, the understanding of the advantage promoted by squeezed reservoir is still an open question. While our results do not directly demonstrate the underlying physical mechanism, it is plausible that their intrinsically quantum features Walls:83 ; Teich:89 —absent in classical thermal reservoirs—modify the effective decay rates and the Liouvillian structure in a way that favors the suppression of specific decay channels, depending on the squeezing parameters. This could, in principle, create more favorable conditions for the observation of the Mpemba effect.
As final remarks, we note that although our analysis centers on a two-level system in a Markovian regime, the framework we developed is general and applicable to more complex quantum systems. The sufficient condition for the quantum Mpemba effect (QMpE), derived from the Liouvillian spectral structure, is not limited by system dimensionality and it is supported by recent experimental observations in higher-dimensional platforms. Moreover, while additional decoherence mechanisms—such as dephasing or coupling to uncontrolled environmental modes—may shift the regime where the effect is most prominent, they do not preclude its occurrence, as it is governed by population relaxation pathways that remain active even under partial loss of coherence. The device-independent nature of our criteria ensures their robustness across a broad class of dissipative models. Extending this framework to non-Markovian dynamics, particularly involving squeezed or structured reservoirs, presents a promising avenue for further exploration of coherence, dissipation, and thermalization in quantum systems.